Comerpickles:

What is the equation of a line that passes through the point (9, −3) and is parallel to the line whose equation is 2x−3y=6 ?

Comerpickles:

@Vocaloid

Vocaloid:

first find the slope of 2x−3y=6 by re-writing in y = mx + b form

Comerpickles:

m=2/3?

Vocaloid:

well done now, we use that slope to find the equation of a line w/ slope (2/3) and passes through (9,-3) so we can plug these into y = mx + b to find b y = mx + b -3 = (2/3)(9) + b solve for b.

Comerpickles:

How do I do that again?

Vocaloid:

first, let's simplify (2/3)(9) to get 6 now we have -3 = 6 + b then subtract 6 from both sides to get b = -9

Vocaloid:

then putting everything together we get y = (2/3)x - 9 = your answer

Comerpickles:

Written like this? http://prntscr.com/hvnk7s

Comerpickles:

Hello?